OpenStudy (anonymous):

limit as x approaches 0 of: (tan(x)-sin(x))/sin^3(x) I've only thought of simplyfying it to this form: (1-cos(x))/(cos(x)*sin^2(x))

OpenStudy (anonymous):

\[\frac{ 1-\cos x }{ \cos x * \sin^2x }\]

OpenStudy (anonymous):

Well, I got to \[\frac{ \cos x - 1 }{ \sin^2 x } + \frac{ 1 }{ \cos x }\]

OpenStudy (anonymous):

I've done it :) \[\frac{\cos x−1}{\sin^2x}+\frac{1}{\cos x}=\] \[= \frac{1}{\cos x+1} +\frac{1}{\cos x} \]

OpenStudy (anonymous):

I mean \[-\frac{ 1 }{ \cos x + 1 } + \frac{1}{\cos x}\]